OpenStudy (anonymous):

0.05x^2 -x+1 = -1+1 Im trying to figure out what would be my next step to finding the solutions through Completeing the square

OpenStudy (anonymous):

0.05x^2 -x+1 = -1+1 0.05x^2 -x+1 = 0 x^2 -20x+20 = 0 then using quadratic formula do you know??

OpenStudy (anonymous):

how did you get that third step?

OpenStudy (anonymous):

ok x = (-b +- sqrt(b^2-4ac)/2a x=(-(-20) +- sqrt((-20)^2 - (4 x1x 20))/2 x = (20 +- sqrt(320))/2 x = (20 +- 17.88)/2 So x = (20 + 17.88)/2 = 18.94 or x = (20 - 17.88)/2 = 1.06

OpenStudy (anonymous):

divided all by 0.05 for third step @FlyinSolo_424

OpenStudy (anonymous):

0.05x^2 -x+1 = -1+1 0.05x^2 -x+1 = 0 ((0.05x^2)/0.05) -(x/0.05)+(1/0.05) = (0/0.05) x^2 -20x+20 = 0

OpenStudy (anonymous):

Thank you So much For explaining I did it once I got that little thing nd I got the same anwser

OpenStudy (anonymous):

welcome @FlyinSolo_424 , good job :)